# Information Gain-Guided Causal Intervention for Autonomous Debiasing Large Language Models

### v1

@@ Abstract
Despite significant progress, recent studies indicate that current large language models (LLMs) may still capture dataset biases and utilize them during inference, leading to the poor generalizability of LLMs. However, due to the diversity of dataset biases and the insufficient nature of bias suppression based on in-context learning, the effectiveness of previous prior knowledge-based debiasing methods and in-context learning based automatic debiasing methods is limited. To address these challenges, we explore the combination of causal mechanisms with information theory and propose an information gain-guided causal intervention debiasing (IGCIDB) framework. This framework first utilizes an information gain-guided causal intervention method to automatically and autonomously balance the distribution of instruction-tuning dataset. Subsequently, it employs a standard supervised fine-tuning process to train LLMs on the debiased dataset. Experimental results show that IGCIDB can effectively debias LLM to improve its generalizability across different tasks.
Information Gain-Guided Causal Intervention for Autonomous Debiasing Large Language Models
Zhouhao Sun1, Xiao Ding1thanks: Corresponding Author, Li Du2, Yunpeng Xu1, Yixuan Ma1, Yang Zhao1, Bing Qin1, Ting Liu1 1Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China 2Beijing Academy of Artificial Intelligence, Beijing, China {zhsun, xding, ypxu, yxma, yzhao, bqin, tliu}@ir.hit.edu.cn duli@baai.ac.cn
@@ 1 Introduction
Large language models (LLMs), through the pre-training and supervised fine-tuning process, have demonstrated remarkable ability to follow human instructions to solve various tasks (Achiam et al., 2023), demonstrating immense potential in real-world applications.
Figure 1: The examples in dataset before and after data debiasing. By intervening example 2, answer with less-known individual are introduced (MORGAN Moses) so that LLMs cannot learn to make predictions solely relying on an individual’s popularity.
@@
Although achieving promising performance, LLMs would also inevitably learn dataset bias during the supervised fine-tuning process, such as negation bias and popularity bias Schick et al. (2021); Mukherjee et al. (2021); Navigli et al. (2023); Klimashevskaia et al. (2024). These dataset biases would lead to poor generalizability of LLMs (Yang et al., 2024; Cheng and Amiri, 2024; Yang et al., 2025). As shown in Figure 1, the original dataset only contains samples where the answers are well-known figures, and it lacks samples featuring less-known individuals as answers. Consequently, LLMs trained on such instruction-tuning corpus would also learn to utilize the biased feature ’popularity of the individuals’ for making predictions. However, when less-known people are generally the correct answer in some datasets, the performance of the LLM will significantly decline, indicating that dataset biases would reduce the generalizability of LLMs.
@@
This problem highlights the necessity of debiasing LLMs. While contemporary debiasing methods are relatively effective, most of them require prior knowledge of dataset biases (Oba et al., 2023; Liu et al., 2023; Cheng and Amiri, 2024). However, dataset biases can vary significantly across different datasets and tasks (Poliak et al., 2018; Schuster et al., 2019; Schick et al., 2021), making it impractical to identify them one by one manually. Furthermore, a vast amount of biases still remain unrecognized in different datasets (Nie et al., 2020) and new biases are constantly being discovered.
@@
Hence, there is an urgent demand for methods to automatically debias LLMs without prior knowledge of dataset biases. However, previous automatic debiasing approaches (Sun et al., 2024) for LLMs are based on in-context learning (ICL) which can only suppress bias in LLMs instead of completely eliminating it in theory, as they depend on the instruction to mitigate bias without altering the model’s intrinsic parameters.
@@
To address these issues, we start from the essential cause to debias the instruction-tuning dataset and propose a Information Gain-guided Causal Intervention DeBiasing Framework (IGCIDB). To eliminate biased features within the instruction-tuning dataset, it is essential to ensure that these features do not provide any additional information to predict the answers, i.e., the information gain (Csiszár and Körner, 2015) of these features for predicting the answers needs to be 0. Therefore, IGCIDB initially employs the CAL Sun et al. (2024) method to automatically detect biased features within the datasets. Subsequently, the information gain is utilized to check the relationships between biased features and answers. Following this, IGCIDB utilize a causal intervention-based data rewriting method to intervene in the biased features. This process can modify the joint distribution between biased features and answers to eliminate their correlations, effectively reducing the information gain. Finally, LLMs are fine-tuned on the debiased dataset utilizing the supervised fine-tuning method.
@@
As illustrated in Figure 1, the original dataset is biased because it exclusively contains data whose answer is a well-known individual, making LLMs trained on such dataset rely sorely on the individual’s popularity when answering similar questions. By utilizing the information gain-guided causal intervention method to revise example 2, the proportion of well-known individual (James) and less-known individual (MORGAN Moses) in the answer becomes the same. Consequently, the biased feature individual’s popularity does not provide any information gain to predict the answers which indicates the success of the data debiasing process.
@@
Experimental results show that our approach can improve the generalizability of LLMs on transfer and challenge datasets for different tasks. We also note that dataset debiased by a certain LLM can also be utilized for another LLM to mitigate bias.
@@ 2 Biased Features within the Instruction-tuning Datasets
Text records and reflects human thoughts. However, due to inherent prejudice or preference of human beings, as well as potential annotation artifacts, various biases such as gender and popularity biases still broadly exist in various datasets Steen and Markert (2024); Dai et al. (2024).
@@
In instruction-tuning datasets, each piece of data consists of a question (i.e., instruction) and an answer. And each instruction-tuning dataset contains three types of features: The first category of features provides no assistance in accurately answering the questions within the dataset and is referred to as the non-predictive feature. The remaining features are collectively termed predictive features. Among these predictive features, causal features directly determine the answer and exhibit a causal relationship with it. The other predictive features, which do not have a direct causal relation to the answer, can be defined as predictive but non-causal features. For instance, in many datasets, the presence of negation word such as ‘not’ often appears when the sentiment of the sentence is negative rather than positive (Mukherjee et al., 2021). However, the mere presence of the word ‘not’ does not necessarily indicate that the sentiment is negative. In this case, the presence of the word "not" serves as a predictive but non-causal feature. In the debiasing scenario, the predictive but non-causal feature are also referred to as biased features.
Among these categories of features, non-predictive features are generally not learned and utilized by powerful LLMs to answer questions, as they do not contribute to providing correct answers. Predictive features, which are useful for accurately answering questions within the dataset, are typically learned by LLMs. However, biased features lack a causal relationship with the answers. Considering that data distribution varies across different datasets, the predictive power of these biased features for answers may change. For example, if the sentiment of sentences in a dataset is generally positive when the negation word ‘not’ appears, the performance of LLMs have learned to predict negative sentiment based on the negation word may decline. Consequently, compared to learning causal features, the out-of-distribution generalization ability of LLMs may be decreased by learning and utilizing biased features.
@@ 3 Method
As illustrated in Figure 2 (a), IGCIDB contains three main steps: (i) Fine-tune the pre-trained LLM on original biased instruction-tuning dataset; (ii) Biased feature identification and information gain-guided causal intervention for data debiasing. (iii) Fine-tune the pre-trained LLM on unbiased instruction-tuning dataset. Since (i) and (iii) are standard supervised fine-tuning processes, our focus will be on providing a detailed explanation on process (ii).
@@ 3.1 Biased Feature Identification
For brevity, we denote the LLM fine-tuned on the original biased instruction-tuning datasets as the biased LLM. To automatically identify explainable biased features, we utilize the causal-guided active learning (CAL) method proposed by Sun et al. (2024). This method is applied independently to each task within the original instruction-tuning dataset. Through this approach, we are able to identify a series of biased features using the biased LLM. Though there are certain limitations, such as being unable to deal with open-ended tasks without standard answers (e.g. writing essays), it is currently the only automated method for identifying interpretable biased features.
Specifically, for discriminative tasks, we directly apply the original CAL method to automatically identify biased features. For other tasks with standard answers such as the QA task, we apply the CAL method to data with consistent question objectives [^1] The consistency of question objectives also means that the answer space of these data to be the same for each category of data to identify biased features, since the CAL method requires uniformity in the question’s goal.
@@ 3.2 Information Gain-guided Data Debiasing Goal
To debias the instruction-tuning dataset, we need to ensure that the identified biased features do not provide any additional information to predict the answers. That is, we need to make the information gain (Csiszár and Körner, 2015) of these features for predicting the answers to be 0 under this dataset. Specifically, for any biased feature $B$:
@@
where $IG$ is the information gain, and $Y$ is the answer. And it can be satisfied if and only if:
@@
The specific derivation process for Eq. 2 can be found in the Appendix A. Since there is no prior knowledge about $Y$, the probability distribution of $P(Y)$ should be uniform to avoid introducing label bias. Consequently, for any two values $y_{i}$ and $y_{j}$ of $Y$, we have
@@
where $N(y_{i},b_{k})$ denotes the number of instances whose answer and biased feature are $y_{i}$ and $b_{k}$, respectively. The information gain-guided data debiasing goal is shown in Figure 2 (b). Note that even if the number $N$ exhibits slight disparities, LLMs are unlikely to learn these subtle correlations. Therefore, we relax the strict equality in Eq. 4 to an approximate equality. To the best of our knowledge, we are the first to establish strict criteria that an unbiased instruction-tuning dataset must satisfy.
For scenarios where the value space of the biased feature $B$ and answer $Y$ can be easily enumerated, such as the negation bias [^2] A negation word either exists in the sentence or not. in the sentiment analysis task, this data debiasing goal can directly guide the subsequent causal intervention process. However, for scenarios where the value space of the biased feature $B$ or the answer $Y$ is difficult to enumerate, the data debiasing goal should be adjusted. For example, if the biased feature is the individual’s popularity in the given context and the question is who caused a certain event according to the given context, the value space of the biased feature will be the popularity of people appearing in the context, and the value space of the answer will be the people appearing in the context. Since the people appearing in each context are different, it is difficult to enumerate them one by one. Additionally, it does not matter who the people are appearing in the context, but the popularity of the people matters. As a result, the value space of the biased feature can be changed to the ratio of popular people and unpopular people appearing in the context, and the value space of the answer can be changed to the popularity of the answer character. Other data debiasing goals that need to be adjusted and the adjustment process can be found in the Appendix B.
@@ 3.3 Causal Intervention-based Data Rewriting
The process of causal intervention (Pearl, 2009) involves modifying the original data distribution to produce a distribution that meets our expectations, which enables the realization of the information gain-guided data debiasing goal. Specifically, we first analyze the original distribution of the dataset and determine the intervention direction (e.g., the arrow in Figure 2 (c)). Subsequently, a causal intervention procedure is applied to a proportion of data by executing the operation $do(B=b_{k})$, thereby modifying the data distribution to meet the data debiasing goal. This causal intervention process intervenes in the biased features, without changing the other semantics of the samples [^3] For semantic that tightly tied to biased features, we also need to modify the semantics of the sample. For example, the individuals are tied to popularity, so we need to modify the figure when modifying popularity..
@@
To perform the operation $do(B=b_{k})$, we utilize the LLM itself to rewrite the data. Specifically, this causal intervention process employs the few-shot prompting method to rewrite samples within the dataset using the LLM itself. To ensure that the biased feature of the rewritten samples satisfies the intervention objectives $b_{k}$, we use the LLM itself or some metrics to check different biased features (specific details can be found in Appendix C). If the requirements are met, the rewritten data is retained. Otherwise, the rewriting process is repeated until the intervenion objectives are met or the rewriting fails to meet the objectives three times.
NLI PI SA QA
Llama3.1 MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla SFT 83.5 72.7 86.6 75.0 95.4 95.4 87.3 65.6 86.5
RAZOR 83.7 73.2 86.4 74.7 95.3 95.6 87.3 65.6 86.6
IGCIDB 83.4 75.4 86.2 77.3 95.4 95.5 87.6 66.2 86.8
Gemma2 MNLI SNLI QQP MRPC SST2 IMDB Squadv1 TriviaQA Squadv2
Vanilla SFT 79.5 65.7 86.8 73.8 95.9 95.3 85.5 68.8 84.8
RAZOR 79.4 67.6 86.7 72.9 95.8 95.5 85.5 68.9 84.9
IGCIDB 79.1 69.0 86.5 77.6 95.8 95.5 85.6 69.1 85.2
Table 1: Comparison of IGCIDB with baselines across two LLMs on ID and transfer test sets.
Note that the number of samples used for the causal intervention process can be automatically calculated. For simplicity, we manually determine the intervention direction and the number of samples to be intervened in order to satisfy the data debiasing goal.
@@ 4.1 Experimental Details
In this work, we use Llama3.1-8B Dubey et al. (2024) to debias the instruction-tuning dataset, and then fine-tune Llama3.1-8B on the debiased dataset for experiments. To examine the generalizablity of the debiased dataset, Gemma2-9B Team et al. (2024) is also utilized to fine-tune upon the debiased dataset.
Our instruction-tuning dataset consists of Lima (Zhou et al., 2024) and a portion of Flan 2021 (Wei et al., 2022) datasets. Due to the time limit, we separately choose 10,000 random samples from 5 sub-datasets of the Flan 2021 dataset, including MNLI (Williams et al., 2018), QQP (Sharma et al., 2019), SST2 (Socher et al., 2013), Squadv1 (Rajpurkar, 2016), and TriviaQA (Joshi et al., 2017). These sub-datasets include four types of tasks: natural language inference (NLI), paraphrase identification (PI), sentiment analysis (SA), and question answer (QA). Since the task of the Lima dataset is multi-turn dialogue which is an open-ended task without standard answers, the biased features cannot be automatically induced by the CAL method Sun et al. (2024), and hence we apply the IGCIDB method to the other four tasks.
We utilized lora (Hu et al., 2022) for supervised fine-tuning (SFT) in our experiments. During the training process, we employed a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 1e-4 for both models. We do not use early stopping, but instead train for three epochs. More details can be seen in Appendix D.

### v2

@@ Abstract
Despite significant progress, recent studies indicate that current large language models (LLMs) may still capture dataset biases and utilize them during inference, leading to the poor generalizability of LLMs. However, due to the diversity of dataset biases and the insufficient nature of bias suppression based on in-context learning, the effectiveness of previous prior knowledge-based debiasing methods and in-context learning based automatic debiasing methods is limited. To address these challenges, we explore the combination of causal mechanisms with information theory and propose an information gain-guided causal intervention debiasing (ICD) framework. To eliminate biases within the instruction-tuning dataset, it is essential to ensure that these biases do not provide any additional information to predict the answers, i.e., the information gain of these biases for predicting the answers needs to be 0. Under this guidance, this framework utilizes a causal intervention-based data rewriting method to automatically and autonomously balance the distribution of instruction-tuning dataset for reducing the information gain. Subsequently, it employs a standard supervised fine-tuning process to train LLMs on the debiased dataset. Experimental results show that ICD can effectively debias LLM to improve its generalizability across different tasks.
@@ 1 Introduction
Large language models (LLMs), through the pre-training and supervised fine-tuning process, have demonstrated remarkable ability to follow human instructions to solve various tasks [1], demonstrating immense potential in real-world applications.
@@
Although achieving promising performance, LLMs would also inevitably learn dataset bias during the supervised fine-tuning process, such as negation bias and popularity bias [30, 21, 22, 13]. These dataset biases would lead to poor generalizability of LLMs [46, 3, 47]. As shown in Figure 1, the original dataset only contains samples where the answers are well-known figures, and it lacks samples featuring less-known individuals as answers. Consequently, LLMs trained on such instruction-tuning corpus would also learn to utilize the biased feature ‘popularity of the individuals’ for making predictions. However, when less-known people are generally the correct answer in some datasets, the performance of the LLM will significantly decline, indicating that dataset biases would reduce the generalizability of LLMs.
@@
This problem highlights the necessity of debiasing LLMs. While contemporary debiasing methods are relatively effective, most of them require prior knowledge of dataset biases [24, 16, 3]. However, dataset biases can vary significantly across different datasets and tasks [26, 31, 30], making it impractical to identify them one by one manually. Furthermore, a vast amount of biases still remain unrecognized in different datasets [23] and new biases are constantly being discovered.
Figure 1: The examples in dataset before and after data debiasing. By intervening example 2, an answer with a less-known individual is introduced (MORGAN Moses) so that LLMs cannot learn to make predictions solely relying on an individual’s popularity.
@@
Hence, there is an urgent demand for methods to automatically debias LLMs without prior knowledge of dataset biases. However, previous automatic debiasing approaches [35] for LLMs rely on in-context learning (ICL) which can only suppress bias in LLMs instead of completely eliminating it in theory, as they mitigate bias through instructions without altering the model’s intrinsic parameters.
@@
To address these issues, we start from the essential cause to debias the instruction-tuning dataset and propose an Information gain-guided Causal intervention Debiasing framework (ICD). To eliminate biased features within the instruction-tuning dataset, it is essential to ensure that these features do not provide any additional information to predict the answers, i.e., the information gain [4] of these features for predicting the answers needs to be 0. Therefore, ICD initially employs an automatic bias detection method to identify biased features within the datasets. Subsequently, the information gain is utilized to check the relationships between biased features and answers. Following this, ICD utilize a causal intervention-based data rewriting method to intervene in the biased features. This process can modify the joint distribution between biased features and answers to eliminate their correlations, effectively reducing the information gain. Finally, LLMs are fine-tuned on the debiased dataset utilizing the supervised fine-tuning method.
@@
As illustrated in Figure 1, by utilizing information gain-guided causal intervention method to revise example 2, the proportion of well-known individual (James) and less-known individual (MORGAN Moses) in the answer becomes the same. Consequently, the biased feature individual’s popularity provide no information gain to predict the answers, indicating the success of the debiasing process.
@@
Experimental results show that our approach can improve the generalizability of LLMs on transfer and challenge datasets for four tasks. Furthermore, experiments on three general ability benchmarks demonstrate that IDB can improve the general ability of LLMs. We also note that datasets debiased by a certain LLM can also be utilized for another LLM to mitigate biases.
@@ 2 Biased Features within the Instruction-tuning Datasets
Text records and reflects human thoughts. However, due to inherent prejudice or preference of human beings, as well as potential annotation artifacts, various biases such as gender and popularity biases still broadly exist in various datasets [34, 5].
@@
In instruction-tuning datasets, each piece of data consists of a question (i.e., instruction) and an answer. And each instruction-tuning dataset contains three types of features: The first category of features provides no assistance in accurately answering the questions within the dataset and is referred to as the non-predictive feature. The remaining features are collectively termed predictive features. Among these predictive features, causal features directly determine the answer and exhibit a causal relationship with it. The other predictive features, which do not have a direct causal relation to the answer, can be defined as predictive but non-causal features. For instance, in many datasets, the presence of a negation word such as ‘not’ often appears when the sentiment of the sentence is negative rather than positive [21]. However, the mere presence of the word ‘not’ does not necessarily indicate that the sentiment is negative. In this case, the presence of the word "not" serves as a predictive but non-causal feature. In the debiasing scenario, the predictive but non-causal features are also referred to as biased features.
Among these categories of features, non-predictive features are generally not learned and utilized by powerful LLMs to answer questions, as they do not contribute to providing correct answers. Predictive features, which are useful for accurately answering questions within the dataset, are typically learned by LLMs. However, biased features lack a causal relationship with the answers. Considering that data distribution varies across different datasets, the predictive power of these biased features for answers may change. For example, if the sentiment of sentences in a dataset is generally positive when the negation word ‘not’ appears, the performance of LLMs have learned to predict negative sentiment based on the negation word may decline. Consequently, compared to learning causal features, the out-of-distribution generalization ability of LLMs may be decreased by learning biased features.
@@ 3 Method
As illustrated in Figure 2 (a), ICD contains three main steps: (i) Fine-tune the pre-trained LLM on original biased instruction-tuning dataset; (ii) Biased feature identification and information gain-guided causal intervention for data debiasing. (iii) Fine-tune the pre-trained LLM on unbiased instruction-tuning dataset. Since (i) and (iii) are standard supervised fine-tuning processes, our focus will be on providing a detailed explanation on process (ii).
@@ 3.1 Biased Feature Identification
For brevity, we denote the large language models fine-tuned on the original biased instruction-tuning datasets as the biased LLM. This process aims at automatically identifying explainable biased features within the biased LLM without human prior knowledge. Specifically, we utilize the causal-guided active learning (CAL) method Sun et al. [35], which is currently the only automatic method for identifying interpretable biased features. This method is applied independently to each task within the original instruction-tuning dataset to identify a series of biased features using the biased LLM. The identified biased features for each task can be seen in the Appendix G.
@@ 3.2 Information Gain-guided Data Debiasing Goal
To debias the instruction-tuning dataset, we need to ensure that the identified biased features do not provide any additional information to predict the answers. That is, we need to make the information gain [4] of these features for predicting the answers to be 0 under this dataset. Specifically, for any biased feature $B$:
@@
where $IG$ is the information gain, and $Y$ is the answer. From Eq. 1, we can know that:
$H(Y|B)=H(Y).$ (2)
where H is the information entropy. By expressing $H(Y|B)$ and $H(Y)$ using probability formulas and extracting the public item, we have:
$\sum_{b\in B}\sum_{y\in Y}P(b)P(y|b)(\log_{2}P(y)-\log_{2}P(y|b))=0.$ (3)
The specific derivation process for Eq. 3 can be found in Appendix B. By ensuring that $P(y)=P(y|b)$ for any values of y and b, Eq. 3 can hold true constantly, that is to say:
@@
Since there is no prior knowledge about $Y$, the probability distribution of $P(Y)$ should be uniform to avoid introducing label bias. Consequently, for any two values $y_{i}$ and $y_{j}$ of $Y$, we have:
@@
where $N(y_{i},b_{k})$ denotes the number of instances whose answer and biased feature are $y_{i}$ and $b_{k}$, respectively. The Eq. 6 formalizes the information gain-guided data debiasing goal. To the best of our knowledge, we are the first to establish strict criteria that an unbiased instruction-tuning dataset must satisfy. Note that even if the number $N$ exhibits slight disparities, LLMs are unlikely to learn these subtle correlations. Therefore, we relax the strict equality in Eq. 6 to an approximate equality as shown in Figure 2 (b).
For scenarios where the value space of the biased feature $B$ and answer $Y$ can be easily enumerated, such as the negation bias in the sentiment analysis task [^1] A negation word either exists in the sentence or not, and the value space of the answer for sentiment analysis task is positive and negative., this data debiasing goal can directly guide the subsequent causal intervention process. However, for scenarios where the value space of the biased feature $B$ or the answer $Y$ is difficult to enumerate, we will utilize an automatic regularization method. The automatic regularization method can be divided into two aspects: On the one hand, for the biased feature whose value space is a continuous space, it is normalized to a smaller, limited discrete space. For example, the value space of a person’s popularity is originally a continuous space, but it can be directly normalized into "low popularity" and "high popularity" (this process can be automatically achieved by using LLMs). On the other hand, for answers whose value space is an extensive value space, automatic regularization can be executed based on the biased feature. Take the QA task and the individual’s popularity as the biased feature as an example once more. The value space of the answer can potentially cover any person, so it has a vast value space. Hence, we can directly regularize the person according to the popularity (high or low). This process is also reasonable because when the biased feature under consideration is the individual’s popularity, it does not matter who the people are appearing in the answer, but the popularity of the people matters. Other data debiasing goals that need to be regularized and the specific regularization process can be found in the Appendix C.
@@ 3.3 Causal Intervention-based Data Rewriting
The process of causal intervention [25] involves modifying the original data distribution to produce a distribution that meets our expectations, which enables the realization of the information gain-guided data debiasing goal. Specifically, we first analyze the original distribution of the dataset and determine the intervention direction (e.g., the arrow in Figure 2 (c)). Subsequently, a causal intervention procedure is applied to a proportion of data by executing the operation $do(B=b_{k})$, thereby modifying the data distribution to meet the data debiasing goal. This causal intervention process intervenes in the biased features, without changing the other semantics of the samples [^2] For semantic that tightly tied to biased features, we also need to modify the semantics of the sample. For example, the individuals are tied to popularity, so we need to modify the figure when modifying popularity..
@@
To perform the operation $do(B=b_{k})$, we utilize the LLM itself to rewrite the data. Specifically, this causal intervention process employs the few-shot prompting method to rewrite samples within the dataset using the LLM itself. To ensure that the biased feature of the rewritten samples satisfies the intervention objectives $b_{k}$, we use the LLM itself or some metrics to check different biased features (specific details can be found in Appendix D). If the requirements are met, the rewritten data is retained. Otherwise, the rewriting process is repeated until the intervenion objectives are met or the rewriting fails to meet the objectives three times.
Note that we also hope to minimize the amount of data to be rewritten, so the process of determining the intervention direction and the number of samples to be intervened can be regarded as a linear programming problem. Its objective function is to minimize the amount of rewritten data, and the constraint is that the distribution of the rewritten dataset should satisfy the information gain-guided data debiasing goal. By solving this linear programming problem, we can obtain the intervention direction and the number of samples to be intervened.
@@ 4.1 Experimental Details
In this work, we use Llama3.1-8B [8] to debias the instruction-tuning dataset, and then fine-tune Llama3.1-8B on the debiased dataset for evaluating the ICD method. To examine the generalizablity of the debiased dataset, Gemma2-9B [38] is also utilized to fine-tune upon the debiased dataset.
Our instruction-tuning dataset consists of Lima [49] and a portion of Flan 2021 [43] datasets. Due to the time limit, we separately choose 10,000 random samples from 5 sub-datasets of the Flan 2021 dataset, including MNLI [44], QQP [32], SST2 [33], Squadv1 [27], and TriviaQA [12]. These sub-datasets include four types of tasks: natural language inference (NLI), paraphrase identification (PI), sentiment analysis (SA), and question answer (QA). For NLI, SA and PI task, we ensure that the label is balanced when sampling. Since the task of the Lima dataset is multi-turn dialogue which is an open-ended task without standard answers, the biased features cannot be automatically induced by the CAL method [35], and hence we apply the ICD method to the other four tasks.
We utilized few-shot prompt method for data rewriting and lora [11] for supervised fine-tuning (SFT) in our experiments. During the training process, we employed a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 1e-4 for both models. We do not use early stopping, and train for three epochs instead. And each result is the average and the standard deviation of the scores across three different runs. Llama3.1-8B and Gemma2-9B are trained on 1 NVIDIA A100 80GB PCIe GPU for 3 and 4 hours, respectively. The specific prompt for data rewriting can be seen in the supplemental material.